ADNAN MUHAMMED

Grade 12,

A convex lens of focal length 0.5m and concave lens of focal length 1m are combined. The power of resulting lens will be

A convex lens of focal length 0.5m and concave lens of focal length 1m are combined. The power of resulting lens will be

Grade:12th pass

1 Answers

Khimraj
7 years ago
Resultant power of combination of two lens is given by P = P1+P2So the resultant power is P = 1/f1 + 1/f2P = 1/0.5 + 1/(-1) = 2 DPlease if you like my answer then approve the answer